Trigger Operation Meeting (1/29/08)
Attendees: Rainer Bartoldus, Su Dong, Mark Tibbetts, Debbie Bard, Kim Hojeong, Steve Sekula.
Date : January 29th, 2008
Time : 2:00pm - 3:30pm PST
SLAC Room : Madrone conf room, ROB
Dial-in : 1-510-665-5437, pass code 6649#
--- Minutes ---
- Run Coordinators want to know when pi0pi0 filter will be ready to be implimented. They want to decide when to take off-peak data.
- Conclusion: since we will take off-peak data at the end of the run again, we'll take off-peak data with current trigger configurations. We have to emulate last two trigger configurations already, don't want to add another different configuration.
- Calibration failed. Updated reference file. Calibration is fine.
Trickle injection monitoring
- L1 rate at ~7000 Hz is a clear wall (dead time vs. L1 rate for a week, as of today).
- We'll apply tigher cuts to reduce L3 rates (which is reaching a limit) including SingleGammaFilters with angular cuts.
- L1 trigger modification plan: 2M1G and M*1G0B will replace two 1zxx lines.
- All these changes will be in after off-peak runs. (The off-peak runs will take about 1.5 days.)
- Rainer installed a new Level 3 EPICS server (L3TPcaServer) that provides hadronic A and B ratios and their errors sampled over 10s, 1min and 5min intervals. The channel names are L3T:HADA:1MIN and similar. Sunday Day shift.
This page was created by Kim Hojeong; however, you are welcome to edit and/or correct above items, specially for those who attended the meeting. Thanks.
Last Modified: January 29th, 2008.